Could you please help me with these problems?
1. assume variables are positive and multiply
y^6/9 ( y^-6/9 + 4y ^3/9)
2. Assume x, y, z are positive #'s use properties of logarithms to write expression log
7 sqrt x^6 y ^8 / z^7
3. Find value of x: logx 121/9 = 2
4. simplify (256 x^4 y) 1/4 / (81xy^3) 3/4

Answer by stanbon(75887)

1. assume variables are positive and multiply
y^6/9 ( y^-6/9 + 4y ^3/9)
=y^0 + 4y^1
=1+4y
========================
2. Assume x, y, z are positive #'s use properties of logarithms to write expression log 7 sqrt(x^6 y ^8) / z^7
Comment: Keep in mind this is all base 7.
= log [(x^3y^4)/z^7]
= 3logx + 4logy - 7logz
========================

3. Find value of x: logx 121/9 = 2
x^2 = 121/9
x = 11/3
=============

4. simplify (256 x^4 y)^(1/4) / (81xy^3)^(3/4)
= 4xy^(1/4) / 27x^(3/4)y^(9/4)
= [4/27][x^(1/4)] / y^2
=[4x^(1/4)]/[27y^2]
=================
Cheers,
Stan H.